# Fréchet Power-Scenario Distance: A Metric for Evaluating Generative AI Models across Multiple Time-Scales in Smart Grids

### v1

@@ Abstract
Generative artificial intelligence (AI) models in smart grids have advanced significantly in recent years due to their ability to generate large amounts of synthetic data, which would otherwise be difficult to obtain in the real world due to confidentiality constraints. A key challenge in utilizing such synthetic data is how to assess the data quality produced from such generative models. Traditional Euclidean distance-based metrics only reflect pair-wise relations between two individual samples, and could fail in evaluating quality differences between groups of synthetic datasets. In this work, we propose a novel metric based on the Fréchet Distance (FD) estimated between two datasets in a learned feature space. The proposed method evaluates the quality of generation from a distributional perspective. Empirical results demonstrate the superiority of the proposed metric across timescales and models, enhancing the reliability of data-driven decision-making in smart grid operations.
Index Terms:
Generative model, synthetic data evaluation, machine learning, energy data
@@ I Introduction
As an alternative, distribution-based comparison methods, such as Fréchet Distance (FD), which can capture the broader distributional characteristics of generative models, have been explored in previous work [7]. However, applying FD directly at the data level may not effectively capture the similarity between datasets. Other approaches have opted for task-specific metrics that involve statistical analysis of task-based parameters. For instance, the number of zeros is employed as a task-specific indicator to assess the quality of voltage sag generation in [8]. Similarly, the indicators used to detect anomalies in electricity consumption include residential stimulus allocation, residential tariff allocation, and SME (Small and Medium-sized Enterprise) allocation in [9]. However, these approaches introduce inconsistencies, as the evaluation criteria vary across tasks. The variability makes it challenging to achieve a unified understanding of a model’s overall performance or to establish standardized benchmarks, potentially leading to biased assessments.
@@
Building on insights from computer vision research, where generative model evaluation has been extensively explored, we investigate methods that could establish more reliable evaluation standards. One such approach is the Inception Score (IS) [10], which was initially adopted to assess the quality and diversity of generated images using a pre-trained Inception network. However, IS faced criticism for its inability to detect the mode collapse of the generative model, where the model produces a limited variety of outputs or concentrates on only a few modes within the distribution, resulting in potential failures across various scenarios [11]. The Fréchet Inception Distance (FID) was proposed to overcome these limitations. FID computes the Fréchet distance between the distributions of real and generated images in a vectorized feature space mapped by a pre-trained Inception network, namely Inception v3. This metric considers both the mean and covariance of the feature space of real and generated data distributions, making it more sensitive to variations in both quality and diversity while closely aligning with human perception [12]. [13] suggests using Kernel Inception Distance (KID), which implements Maximum Mean Discrepancy (MMD) with kernel functions applied to the feature space of the Inception model can further improve the metric performance compared to FID. Unlike FID, which can suffer from estimator bias. KID provides an unbiased estimate, ensuring stable evaluation, particularly for small sample datasets. However, this approach is highly dependent on the choice of kernel functions, with different kernels potentially leading to inconsistent results.
Inspired by the success of the FID metric, in this paper we aim to develop a robust metric that could work in the context of multi-time-scale generative data sets in smart grids. The challenge that arises in adapting FID for this context are as follows: (1) The original FID feature extractor, Inception v3, is optimized primarily to capture image-specific features, making it potentially unsuitable for power system data; (2) FID metric is developed based on the ImageNet dataset, which contains over 14 million images across 1,000 categories as a comprehensive benchmark for most image generation tasks [14], whereas there is no standardized dataset of comparable scope and diversity for power systems; (3) Inception V3 model requires a fixed input resolution of 299x299 pixels, which is incompatible with the multi-resolution and multi-duration characteristics inherent to power system data.
To address these three challenges, we (1) train a new feature extraction model to capture both spatial and temporal features critical for power system data, enhancing its performance on such datasets; (2) collect comprehensive data set which, to the best of our knowledge, captures the major synthetic tasks associated with power systems; and (3) design a hierarchical structure composed of multiple feature extractors to handle multi-resolution datasets, enabling features from shorter timescales to be effectively integrated into models operating at longer timescales.
The key contributions of this paper are suggested as follows:
@@
    A hierarchy multi-resolution feature extraction network: We present a novel network design where multiple feature extractors are trained to handle data at different time resolutions and are connected in a hierarchical structure. This approach enables data of varying resolutions to be accommodated within the model, ensuring that features from shorter timescales are effectively transferred to longer timescales, thereby enhancing the model’s ability to capture the complex temporal dynamics inherent in power systems.
@@
    A task-independent metric for generative models in smart grid: To our best knowledge, this is first-of-its-kind introduction of the Fréchet Power-Scenario Distance (FPD) in the smart grids context. FPD is specifically developed and trained to assess the performance of generative models in the power system domain. It can effectively handle multi-timescale data and provides a unified standard for a diverse range of power system tasks.
The rest of the paper is organized as follows: Section II details the design of the proposed evaluation metrics; Section A introduces model training detail; Section IV demonstrates the effectiveness of FPD across various disturbances and its advantages over traditional metrics; and Section V concludes the paper.
@@ II-A Overview of the feature extraction model
We employ neural networks to extract high-level features capable of capturing complex spatial and temporal relations. Currently, a general feature extractor for this task does not exist in the power system domain. Implementing such a general model is challenging due to the diverse time resolutions and durations inherent in power system data. To address this challenge, we build a unified feature extraction framework that accommodates such diversity. As shown in Fig.1, our proposed framework is a multi-resolution multi-duration model, where a series of feature extractors, each trained for a specific time resolution, are interconnected hierarchically. This design allows multiple extractors to be combined for different data durations, with output features from each higher-resolution extractor feeding into the input of the subsequent lower-resolution extractor. This integration of high-resolution features into lower-resolution modules allows for effective utilization of detailed data in lower-resolution models without significantly increasing dimensional complexity.
@@ II-B Hierarchical feature extraction model
Considering the different timescales underlying the steady-state and transient-state data in power systems, we proposed two types of feature extraction models respectively. Due to the different sampling intervals, there are multiple resolutions and durations in steady-state data. Therefore, we adopt a multi-layer hierarchical architecture. For the transient-state situations, the sample intervals are much simpler, typically a selection of integer multiples of the rated frequency. Hence, we designed a single-layer model for the transient-state case.
@@
[…]
@@
To accomodate steady-state data in different resolutions and durations, we employ a multi-layer structure. Each layer is designed to process data at a specific resolution $r_{s}$ and transform it into a feature space aligned with the next higher resolution $r_{s+1}$. We denote the extractor handling data at resolution $r_{s}$ as $\mathcal{M}_{r_{s}}(x_{r_{s}})$, where $x_{r_{s}}$ represents data from the $r_{s}$ resolution. Each encoding function $\mathcal{M}_{r_{s}}$ takes an input $x_{r_{s}}$ The extractor then outputs a feature representation $z_{r_{s}}$ for the next resolution $r_{s+1}$ model processing
@@
We employ these extractors in different resolutions in a hierarchical structure, where each subsequent extractor builds upon the features aggregated from the stack of extractors in the prior level. A switch mechanism is introduced so that each extractor can adapt to the type of input it receives. Each extractor can select its input from either the features extracted at the prior resolution level, or the raw data at the current resolution level, as demonstrated in Fig.1. For instance, the daily module ($\mathcal{M}_{r_{1}}$) can directly process 24 hourly data points if only hourly data is available, producing hourly features on a daily scale. Alternatively, if only 5-minute data is available, 24 hourly modules ($\mathcal{M}_{r_{0}}$) can be used to transform the 5-minute data for each hour (consisting of 12 data points per hour) into corresponding hourly features. Subsequently, the resulting 24 hourly features are concatenated and utilized as input to the daily module.
@@
It is noteworthy that this structure has the capability to adapt multi-resolution, multi-duration data. The input data can be injected at any designated entry point in the model, depending on the data resolution $r_{s}$. The model will process the data through a series of lower-level extractors to higher-level extractors and finally generate the desired feature representation at the module corresponding to the target scale $r_{e}$.
@@
During training and usage, the model processes the data in the form $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$ given the resolution of data $r_{s}$. Here, $N$ denotes the total number of data points across all sequences. For example, consider 1,000 data samples, each representing daily-scale data recorded at 5-minute intervals. Since one day has $288$ such intervals, the total number of data points is $1{,}000\times 288=288{,}000$ $L_{r_{s}}$ is the number of data points required at resolution $r_{s}$ to form a single data point at the next higher resolution $r_{s+1}$; for example, for 5-min resolution data, $L_{\text{5-min}}=12$. Therefore, to process the entire batch of sequences, the dataset must be divided into segments of size $\frac{N}{L_{r_{s}}}$. Under this setup, $\frac{N}{L_{r_{s}}}$ modules $\mathcal{M}_{r_{s}}$ will each handle a segment of length $L_{r_{s}}$, which aligns with the module design in our structure. Next, we define $D_{r_{s}}$ as the augmented feature dimension associated with resolution $r_{s}$. This dimension includes features extracted from the preceding module, as well as a one-dimensional mean-value derived from the prior-level data. In contrast, the raw time-series data remain one-dimensional. In this case, the higher-level data is viewed as the mean-value of the lower-level data values, and the other lower-level extracted features can be simply zero-padded. This dimension, $D_{r_{s}}$, needs to be chosen properly so that feature space distribution is approximately Gaussian distributed because the subsequent Fréchet Distance calculation assumes that the features follow a Gaussian distribution.
@@
$\mathbf{I}_{r_{s}}=\begin{cases}\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right],&\text{if $z_{r_{s-1}}$ is available}\\ \left[0,x_{r_{s}}\right],&\text{if $x_{r_{s}}$ is available}\\ \end{cases}$
In each resolution, we first determine input $\mathbf{I}_{r_{s}}$ between the feature representation from the prior level and the data at the current level. Due to the dimensional inconsistency between $\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right]$, which has dimension $D_{r_{s}}$, and $x_{r_{s}}$, which has dimension $1$, we need to pad $x_{r_{s}}$, if necessary, to match the dimension $D_{r_{s}}$. This operation ensures that all choices of $\mathbf{I}_{r_{s}}$ have the same dimension and can be adapted by the model. Also, considering the days in different months are inconsistent, for $r=2$, we pad the length to 31 if needed. Then we reshape the $\mathbf{I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$. This operation allows each module at a given resolution to process only a segment of the entire input sequence. Then, by stacking multiple such modules together, we ensure that the entire sequence in $r_{e}$ scale is effectively processed. The reshaped feature vector $z_{r_{s}}$ is passed as input to the extractor at the resolution $r_{s+1}$. We then update the resolution to $r_{s}=r_{s+1}$ and repeat this process sequentially until target resolution $r_{e}$ is reached. Once is $r_{e}$ achieved, we output the resulting feature vectors, which can be used to calculate the FPD score.
The parameters and modules corresponding to each resolution $r$ are listed in Table I. The feature extraction algorithm is outlined in Algorithm 1.
TABLE I: Parameters and modules in steady-state model
$i$ $r_{i}$ $D_{r_{i}}$ $L_{r_{i}}$ $\mathcal{M}_{r_{i}}$
0 5-min 1 12 Hourly Module
1 Hourly 10 24 Daily Module
2 Daily 14 31 Monthly Module
3 Monthly 20 12 Yearly Module
4 Yearly
@@
1:  Input: Duration $r_{e}$, Start resolution $r_{s}$, Batch size $N$, Input data $x_{r_{s}}$, Feature dimension $D_{r_{s}}$, Data Length $L_{r_{s}}$ at resolution $r_{s}$
@@
For the architecture design of a single feature extractor in both models, we utilize residual blocks in the feature extractor as demonstrated in Fig.1. The residual connections have been widely studied and proved to increase neural network depth to enhance representation capabilities. We utilize the conv-1d and batchnorm-1d layers for accommodating the time-series inputs. This Resnet-like feature extractor is a base component for different scenarios and timescale tasks, followed by simple linear heads for different downstream training objectives.
Figure 1: Overview of the hierarchical multi-resolution feature extraction framework, integrating multiple time-scale feature extractors to enhance generative model evaluation in power systems.

### v2

# Fréchet Power-Scenario Distance: A Metric for Evaluating Synthetic Data Across Multiple Time-Scales in Smart Grids

@@ Abstract
Generative artificial intelligence (AI) models in smart grids have advanced significantly in recent years due to their ability to generate large amounts of synthetic data, which would otherwise be difficult to obtain in the real world due to confidentiality constraints. A key challenge in utilizing such synthetic data is how to assess the data quality produced from such generative models. Traditional metrics such as sample-wise Euclidean distance and distributional distances applied directly to raw generated data inadequately reflect higher-order temporal dependencies and cross-temporal relationships between real and synthetic series, and thus struggle to discriminate generative quality. In this work, we propose a novel metric based on the Fréchet Distance (FD) estimated between two datasets in a learned feature space. The proposed method assesses synthetic data quality via distributional comparisons in a feature space derived from a model tailored to the smart grid domain. Empirical results demonstrate the superiority of the proposed metric across downstream tasks and generative models, enhancing the reliability of data-driven decision-making in smart grid operations.
@@ I Introduction
As an alternative, distribution-based approaches—covering proper scoring rules such as the Continuous Ranked Probability Score (CRPS) [7] and Energy Score (ES)[8], and distribution-to-distribution distances such as Kullback–Leibler (KL) divergence [9], Maximum Mean Discrepancy (MMD)[10], and Wasserstein distance [11]—have been explored in prior work to capture the broader distributional characteristics of synthetic datasets. However, when computed directly on raw data, distribution-level metrics tend to underweight higher-order structures such as temporal dependencies, which are crucial in smart grid time-series. Consequently, they can fail to discriminate meaningful changes in generative quality that primarily manifest in the dependence structure [12].
@@
Other approaches have opted for task-specific metrics that involve statistical analysis of task-based parameters. For instance, the number of zeros is employed as a task-specific indicator to assess the quality of voltage sag generation in [13]. Similarly, the indicators used to detect anomalies in electricity consumption include residential stimulus allocation, residential tariff allocation, and SME (Small and Medium-sized Enterprise) allocation in [14]. However, these approaches introduce inconsistencies, as the evaluation criteria vary across tasks. Such variability makes it difficult to form a unified view of overall synthetic data quality or to establish standardized benchmarks, potentially leading to biased assessments.
@@
Building on insights from computer vision research, where generative model evaluation has been extensively explored, we investigate methods that could establish more reliable evaluation standards. One such approach is the Inception Score (IS) [15], which was initially adopted to assess the quality and diversity of generated images using a pretrained Inception network. However, IS faced criticism for its inability to detect the mode collapse of the generative model, where the model produces a limited variety of outputs or concentrates on only a few modes within the distribution, resulting in potential failures across various scenarios [16]. The Fréchet Inception Distance (FID) was proposed to overcome these limitations. FID evaluates distributional similarity by taking the Fréchet (squared 2-Wasserstein) distance between the Gaussian embeddings of real and generated images in feature spaces extracted by a pretrained Inception network. [17]
Inspired by the success of the FID metric, we argue that computing a distributional metric (e.g., Fréchet distance) in a learned feature space yields a more comprehensive and fair evaluation of smart grid synthetic datasets. The central challenge is thus obtaining a valid feature extractor that provides features aligned with the smart grid characteristics. Inception-v3, the encoder used by FID, is optimized for images rather than time-series. Thus, when applied to smart grid data, it underrepresents temporal dynamics. Moreover, its fixed $299\times 299$ input requirement is incompatible with the inherently multi-resolution, multi-duration nature of power-system signals. A natural alternative is to consider time-series encoders explicitly designed to capture temporal dynamics and to be flexible with respect to input length.
Although prior work has produced generic time-series encoders (e.g., TS2Vec [18], TimeMixer [19]), they are not plug-and-play for smart grids. While these encoders can handle variable sequence lengths—typically by padding or truncating to a fixed window—their learned representations often emphasize broad statistical or sequential regularities and miss physics-driven structure. In contrast to generic time-series, smart grid signals exhibit distinctive structure: strong diurnal and seasonal periodicities, hard physical constraints (e.g., ramp limits, nonnegativity, no solar at night), and a multi-resolution character. Capturing these attributes is crucial for downstream tasks such as anomaly detection, disturbance classification, and evaluation of generative scenarios, yet generic pretrained encoders frequently fail to preserve them.
To address this gap, we propose a power-scenario-aware feature extractor that explicitly encodes the temporal and spectral signatures of grid data. Leveraging a multi-resolution hierarchical architecture and trained on large-scale power datasets, the extractor uses a classification head to emphasize local dynamics/pattern regimes and a regression head to capture broader statistical trends. Together, these objectives yield representations that faithfully capture the physical underpinnings of power-system data. Computing a distributional distance (e.g., Fréchet) in this feature space then provides a more insightful, effective, and robust evaluation of power-domain generative models.
The key contributions of this paper are as follows:
@@
    A hierarchical multi-resolution feature extraction network: We present a novel network design where multiple feature extractors are trained to handle data at different time resolutions and are connected in a hierarchical structure. This framework accommodates multi-resolution inputs, propagating features learned at shorter timescales to longer horizons. Each extractor is trained on smart grid datasets using joint classification and regression tasks, encouraging representations that capture source-specific dynamics as well as general statistical structure. Collectively, this improves the model’s capacity to represent the complex temporal behavior inherent in smart grids.
@@
    A feature-based, temporally aware distributional metric for generative models in smart grids: Unlike traditional metrics that operate directly on raw data, we propose a metric that evaluates in a learned feature space: specifically, we compute the Fréchet distance between extracted feature embeddings of real and generated data to quantify generative quality. Extensive experiments show that this approach robustly and consistently handles multi-timescale data and provides a unified standard across diverse smart grid generative tasks. We refer to this metric as the Fréchet Power-Scenario Distance (FPD).
The rest of the paper is organized as follows: Section II details the design of the proposed evaluation metrics; Section III introduces model training details; Section IV demonstrates the effectiveness of FPD across various disturbances and its advantages over traditional metrics; and Section V concludes the paper.
@@ II-A Overview of the feature extraction model
We employ neural networks to extract high-level features capable of capturing complex spatial and temporal relations. Currently, a general feature extractor for this task does not exist in the smart grid domain. Implementing such a general model is challenging due to the diverse time resolutions and durations inherent in smart grid data. To address this challenge, we build a unified feature extraction framework that accommodates such diversity. As shown in Fig. 1, our proposed framework is a multi-resolution, multi-duration model, where a series of feature extractors, each trained for a specific time resolution, are interconnected hierarchically. Every feature extractor during training is paired with a classification head that emphasizes local dynamics and a regression head that captures broader statistical trends, together ensuring that the extracted features faithfully represent the characteristics of smart grid data. This design allows multiple extractors to be combined for different data durations, with output features from each higher-resolution extractor feeding into the input of the subsequent lower-resolution extractor. This integration of high-resolution features into lower-resolution modules enables effective use of fine-grained information while avoiding substantial increases in dimensional complexity. At the same time, it ensures that the output features at each level remain closely aligned with the characteristic patterns of smart grid datasets.
@@ II-B Hierarchical feature extraction model
Considering the different timescales underlying the steady-state and transient-state data in smart grids, we propose two types of feature extraction models respectively. Due to the different sampling intervals, there are multiple resolutions and durations in steady-state data. Therefore, we adopt a multi-layer hierarchical architecture. For the transient-state situations, the sampling intervals are much simpler, typically a selection of integer multiples of the rated frequency. Hence, we design a single-layer model for the transient-state case.
@@
$r_{s}\in\{\text{5-min}\rightarrow\text{hourly}\rightarrow\text{daily}\rightarrow\text{monthly}\rightarrow\text{yearly}\}.$
@@
To accommodate steady-state data in different resolutions and durations, we employ a multi-layer structure. Each layer is designed to process data at a specific resolution $r_{s}$ and transform it into a feature space aligned with the following higher resolution $r_{s+1}$. We denote the extractor handling data at resolution $r_{s}$ as $\mathcal{M}_{r_{s}}(x_{r_{s}})$, where $x_{r_{s}}$ represents data from the $r_{s}$ resolution. Each encoding function $\mathcal{M}_{r_{s}}$ takes an input $x_{r_{s}}$. The extractor then outputs a feature representation $z_{r_{s}}$ for the following resolution $r_{s+1}$ model processing.
@@
We employ these extractors at different resolutions in a hierarchical structure, where each subsequent extractor builds upon the features aggregated from the stack of extractors in the prior level. A switch mechanism is introduced so that each extractor can adapt to the type of input it receives. Each extractor can select its input from either the features extracted at the prior resolution level, or the raw data at the current resolution level, as demonstrated in Fig. 1. For instance, the daily module ($\mathcal{M}_{r_{1}}$) can directly process 24 hourly data points if only hourly data is available, producing hourly features on a daily scale. Alternatively, if only 5-minute data is available, 24 hourly modules ($\mathcal{M}_{r_{0}}$) can be used to transform the 5-minute data for each hour (consisting of 12 data points per hour) into corresponding hourly features. Subsequently, the resulting 24 hourly features are concatenated and utilized as input to the daily module.
@@
It is noteworthy that this structure has the capability to adapt to multi-resolution, multi-duration data. The input data can be injected at any designated entry point in the model, depending on the data resolution $r_{s}$. The model will process the data through a series of lower-level extractors to higher-level extractors and finally generate the desired feature representation at the module corresponding to the target scale $r_{e}$.
@@
During training and usage, the model processes the data in the form $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$ given the resolution of data $r_{s}$. Here, $N$ denotes the total number of data points across all sequences. Consider 1,000 data samples, each representing daily-scale data recorded at 5-minute intervals. Since one day has $288$ such intervals, the total number of data points is $1{,}000\times 288=288{,}000.$ $L_{r_{s}}$ is the number of data points required at resolution $r_{s}$ to form a single data point at the next higher resolution $r_{s+1}$. For 5-min resolution data, $L_{\text{5-min}}=12$. Therefore, to process the entire batch of sequences, the dataset must be divided into segments of size $\frac{N}{L_{r_{s}}}$. Under this setup, $\frac{N}{L_{r_{s}}}$ modules $\mathcal{M}_{r_{s}}$ will each handle a segment of length $L_{r_{s}}$, which aligns with the module design in our structure. Next, we define $D_{r_{s}}$ as the augmented feature dimension associated with resolution $r_{s}$. This dimension includes features extracted from the preceding module, as well as a one-dimensional mean value derived from the prior-level data. In contrast, the raw time-series data remain one-dimensional. In this case, the higher-level data are viewed as the mean value of the lower-level data values, and the other lower-level extracted features can be simply zero-padded. The dimension $D_{r_{s}}$ must be selected carefully to ensure that the FPD score computed in this feature space reliably and smoothly reflects the actual differences between datasets with different disturbance levels. To determine an appropriate dimension, we perform ablation tests by varying the feature dimension that achieves the best performance across different levels of Gaussian noise disturbance. An example ablation result for the hourly module is presented in SI Sec.S1; Fig. S1. The selected dimensions $D_{r_{s}}$ for all modules are summarized in SI Sec.S1; Table S4.
@@
[…]
In each resolution, we first determine input $\mathbf{I}_{r_{s}}$ between the feature representation from the prior level and the data at the current level. Due to the dimensional inconsistency between $\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right]$, which has dimension $D_{r_{s}}$, and $x_{r_{s}}$, which has dimension $1$, we need to pad $x_{r_{s}}$, if necessary, to match the dimension $D_{r_{s}}$. This operation ensures that all choices of $\mathbf{I}_{r_{s}}$ have the same dimension and can be adapted by the model. Also, considering the days in different months are inconsistent, for $r=2$, we pad the length to 31 if needed. We then reshape $\mathbf{I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$. This operation allows each module at a given resolution to process only a segment of the entire input sequence. Then, by stacking multiple such modules together, we ensure that the entire sequence in $r_{e}$ scale is effectively processed. The reshaped feature vector $z_{r_{s}}$ is passed as input to the extractor at the resolution $r_{s+1}$. We then update the resolution to $r_{s}=r_{s+1}$ and repeat this process sequentially until the target resolution $r_{e}$ is reached. Once $r_{e}$ is achieved, we output the resulting feature vectors, which can be used to calculate the FPD score.
The parameters and modules corresponding to each resolution $r$ are listed in SI Sec.S1; Table S4. The feature extraction algorithm is outlined in Algorithm 1.
@@
1:Input: Duration $r_{e}$, Start resolution $r_{s}$, Batch size NN, Input data $x_{r_{s}}$, Feature dimension $D_{r_{s}}$, Data Length $L_{r_{s}}$ at resolution $r_{s}$
@@
For the architecture design of a single feature extractor in both models, we utilize residual blocks in the feature extractor as demonstrated in Fig. 1. The residual connections have been widely studied and proven to increase neural network depth to enhance representation capabilities. We utilize the conv-1d and batchnorm-1d layers for accommodating the time-series inputs. This ResNet-like feature extractor is a base component for different scenarios and timescale tasks, followed by simple linear heads for different downstream training objectives.
Figure 1: Overview of the hierarchical multi-resolution feature extraction framework, integrating multiple time-scale feature extractors to enhance synthetic data evaluation in smart grids.